# Real-Time Image Segmentation via Hybrid Convolutional-Transformer Architecture Search

### v1

@@ V Conclusion
In this paper, we propose a novel HyCTAS framework to search for efficient networks with high-resolution representation and attention. We introduce a light-weight convolution module to reduce the computation cost while preserving high-resolution information and a memory efficient self-attention module to capture long-range dependency. By combing a multi-branch search space and an efficient genetic search algorithm with multiple objectives, our searched architectures not only fuse features from selective branches for efficiency, but also maintain high-resolution representations for performance. Besides, the discovered architecture also provides us with some insights into how to use self-attention when designing new networks. Extensive experiments demonstrate that HyCTAS outperforms previous methods on the semantic segmentation task, while significantly improving the inference speed. Future work will focus on applying this approach to more challenging tasks, such as key-point localization and object detection.

### v2

@@ 6 Conclusion
In this paper, we propose a novel HyCTAS framework to search for efficient networks with high-resolution representation and attention. We introduce a lightweight convolution module to reduce computation costs while preserving high-resolution information and a memory-efficient self-attention module to capture long-range dependencies. By combining a multi-branch search space and an efficient genetic search algorithm with multiple objectives, our searched architectures not only fuse features from selective branches for efficiency but also maintain high-resolution representations for performance. Additionally, the discovered architecture provides insights into how to use self-attention when designing new networks. Extensive experiments demonstrate that HyCTAS outperforms previous methods in both semantic segmentation and panoptic segmentation tasks while significantly improving inference speed. Future work will focus on applying this approach to more challenging tasks, such as key-point localization and object detection.
Acknowledgement
This work is jointly supported by National Key Research and Development Program of China (2016YFB1001000), National Natural Science Foundation of China (61525306, 61633021, 61721004, 61420106015, 61806194), Capital Science and Technology Leading Talent Training Project (Z181100006318030), and Beijing Science and Technology Project (Z181100008918010).
Author Contributions
H. Yu conceived the initial idea, designed the supernet, and conducted the majority of experiments. C. Wan and X. Dai contributed to the multi-objective search algorithm implementation and code maintenance. M. Liu, D. Chen, and B. Xiao provided guidance on memory-efficient self-attention design and revised the manuscript. Y. Huang, Y. Lu, and L. Wang brought expertise in high-resolution networks, coordinated the experimental setup, and offered theoretical insights on multi-branch architectures. All authors reviewed the paper, discussed experiments, and approved the final manuscript.